Chris Sandvig Irrigation, Inc., has summarized the price list from four potential suppliers of an underground control valve. See the table below. Annual usage is 1,100 valves; order cost is $14 per order; and annual inventory holding costs are $7.55 per unit. Vendor A Vendor B Vendor C Vendor D Quantity 1-14 15-24 25-99 100-199 200-399 400+ Price $35.00 34.75 33.55 32.35 31.15 30.75 Quantity 1-24 25-99 100-199 200-399 400+ Price $34.75 34.00 32.80 31.60 30.50 Quantity 1-49 50-149 150-299 300+ Price $34.50 33.75 32.50 31.10 Quantity 1-149 150-299 300+ Price $34.25 33.00 31.00 Which vendor should be selected and what order quantity is best if Sandvig Irrigation wants to minimize total cost? Chris Sandvig Irrigation should order units at a time from V (enter your response as a whole number)
Answer #1

EOQ = sqrt(2*Annual demand*Order cost/Holding cost) = sqrt(2*1100*14/7.55) = 64 (Round off)

Total cost = Annual Demand/EOQ*Order cost + EOQ/2*Holding cost + Price*Annual demand

1) Vendor A: EOQ = 64, Price = 33.55

Total cost = 1100/64*14 + 64/2*7.55 + 1100*33.55 = 37387

2) Vendor A: EOQ = 100, Price = 32.35

Total cost = 1100/100*14 + 100/2*7.55 + 1100*32.35 = 36117

3) Vendor A: EOQ = 200, Price = 31.15

Total cost = 1100/200*14 + 200/2*7.55 + 1100*31.15 = 35097

4) Vendor A: EOQ = 400, Price = 30.75

Total cost = 1100/400*14 + 400/2*7.55 + 1100*30.75 = 35374

5) Vendor B: EOQ = 64, Price = 34

Total cost = 1100/64*14 + 64/2*7.55 + 1100*34 = 37882

6) Vendor B: EOQ = 100, Price = 32.8

Total cost = 1100/100*14 + 100/2*7.55 + 1100*32.8 = 36612

7) Vendor B: EOQ = 200, Price = 31.6

Total cost = 1100/200*14 + 200/2*7.55 + 1100*31.6 = 35592

8) Vendor B: EOQ = 400, Price = 30.5

Total cost = 1100/400*14 + 400/2*7.55 + 1100*30.5 = 35099

9) Vendor C: EOQ = 64, Price = 33.75

Total cost = 1100/64*14 + 64/2*7.55 + 1100*33.75 = 37607

10) Vendor C: EOQ = 150, Price = 32.5

Total cost = 1100/150*14 + 150/2*7.55 + 1100*32.5 = 36419

11) Vendor C: EOQ = 300, Price = 31.1

Total cost = 1100/300*14 + 300/2*7.55 + 1100*31.1 = 35394

12) Vendor C: EOQ = 64, Price = 34.25

Total cost = 1100/64*14 + 64/2*7.55 + 1100*34.25 = 38157

13) Vendor C: EOQ = 150, Price = 33

Total cost = 1100/150*14 + 150/2*7.55 + 1100*33 = 36969

14) Vendor C: EOQ = 300, Price = 31

Total cost = 1100/300*14 + 300/2*7.55 + 1100*31 = 35284

Lowest cost scenario is Case 3:

Vendor A

Order quantity = 200